OpenStudy (anonymous):

Transform the formula for the given variable P = 4s Solve for s

jimthompson5910 (jim_thompson5910):

You're multiplying 's' by 4, so you undo this by doing what?

OpenStudy (anonymous):

s=p/4

OpenStudy (anonymous):

really that all it said like i didnt understand it too

OpenStudy (anonymous):

thank you very much but how do you get it???

jimthompson5910 (jim_thompson5910):

You undo multiplying by 4 by dividing both sides by 4

jimthompson5910 (jim_thompson5910):

So you divide 4s by 4 to get 4s/4 = 1s = s So 4s/4 = s

OpenStudy (anonymous):

really??

jimthompson5910 (jim_thompson5910):

and you divide P by 4 to get P/4

jimthompson5910 (jim_thompson5910):

So \[\Large s = \frac{P}{4}\]

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

sweet thank you so much:)

jimthompson5910 (jim_thompson5910):

you're welcome